# Robust Fine-Tuning from Non-Robust Pretrained Models: Mitigating Suboptimal Transfer With Adversarial Scheduling

### v1

@@ Appendix A Experimental Setup Details
We use train-val-test split from Hua et al. (2024) for Caltech, CUB, Stanford Dogs; and from Heuillet et al. (2025) for Aircraft and Stanford Cars. Training augmentations consist of standard preprocessing methods commonly used for ImageNet and high-resolution images (Marcel & Rodriguez, 2016): random horizontal flips ($p=0.5$), color jitter (brightness, contrast, and saturation set to $0.25$), and random rotations. As done in Robustbench (Croce et al., 2021), images are resized to 224x224 with pixel values in the range $[0,1]$, and data normalization and standardization are directly integrated into the model.
@@ B.2 AutoAttack Results
AutoAttack (Croce & Hein, 2020b) is a stronger and more diverse attack on the models, but is more expensive. We evaluate a few cases (SWIN on {Cars, Aircraft} x $\nicefrac{{4}}{{255}}$, $\nicefrac{{8}}{{255}}$}). Results can be found in Table 5 and Figure 10. Although it takes substantially more time, the results are close to evaluations with APGD.
@@ C.1 Ablation and Sensitivity Analysis
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=8/255$
@@
Table 6: Effect of hyperparameters $T_{1}$ and $T_{2}$. The training dynamics can be found in Figure 11 for $\varepsilon_{g}=\nicefrac{{4}}{{255}}$ and Figure 12 for $\varepsilon_{g}=8/255$. (*) RFT-scheduler reported in main text $\left(T_{1}=12,T_{2}=37\right)$.
@@ C.2 Direct Minimization for Expected Robustness
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=8/255$

### v2

@@ Acknowledgements
Reproducibility Statement
Our study is designed to be fully reproducible. All backbones and datasets are publicly available, with details and references provided in Section 6 and Appendix A, where we also cite the prior work underlying our design choices. Details on the estimation of expected robustness are given in Appendix A.
We provide an anonymized GitHub repository containing the implementation, the results of the hyperparameter optimization, all the data used to generate the paper’s figures and tables, and a script to reproduce them. The repository also includes step-by-step instructions for downloading datasets and pretrained models, creating Python environments, and launching experiments.
Finally, details on compute resources and expected run times are reported in Appendix A.
Link to public github repository: https://github.com/ngnawejonas/EpsilonScheduling
@@ Appendix A Experimental Setup Details
We use train-val-test split from Hua et al. (2024) for Caltech, Cub, Stanford Dogs; and from Heuillet et al. (2025) for Aircraft and Stanford Cars. Training augmentations consist of standard preprocessing methods commonly used for ImageNet and high-resolution images (Marcel and Rodriguez, 2016): random horizontal flips ($p=0.5$), color jitter (brightness, contrast, and saturation set to $0.25$), and random rotations. As done in Robustbench (Croce et al., 2021), images are resized to 224x224 with pixel values in the range $[0,1]$, and data normalization and standardization are directly integrated into the model.
@@ B.2 AutoAttack Results
AutoAttack (Croce and Hein, 2020b) is a stronger and more diverse attack on the models, but is more expensive. We evaluate a few cases (SWIN on {Cars, Aircraft} x $\nicefrac{{4}}{{255}}$, $\nicefrac{{8}}{{255}}$}). Results can be found in Table 6 and Figure 10. Although it takes substantially more time, the results are close to evaluations with APGD.
@@ C.1 Ablation and Sensitivity Analysis
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=\nicefrac{{8}}{{255}}$
@@
Table 7: Effect of hyperparameters $T_{1}$ and $T_{2}$. The training dynamics can be found in Figure 11 for $\varepsilon_{g}=\nicefrac{{4}}{{255}}$ and Figure 12 for $\varepsilon_{g}=\nicefrac{{8}}{{255}}$. (*) RFT-scheduler reported in main text $\left(T_{1}=12,T_{2}=37\right)$.
@@ C.2 Direct Minimization for Expected Robustness
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=\nicefrac{{8}}{{255}}$
@@
Appendix D Statistical Significance
We report paired $t$-test statistics comparing RFT-Fix and RFT-Scheduler at $\epsilon=\nicefrac{{4}}{{255}}$ and $\epsilon=\nicefrac{{8}}{{255}}$ in Table 9. These tests assess whether performance differences between the two strategies are statistically significant across downstream tasks. A paired $t$-test measures whether the mean performance difference between two methods is reliably non-zero; small $p$-values indicate that the observed differences are unlikely to occur by chance.
We also report the averages for each metric per model (Table 10) and per dataset (Table 11).
$\epsilon=\nicefrac{{4}}{{255}}$ $\epsilon=\nicefrac{{8}}{{255}}$
Metric n_pairs $t$-stat $p$-value $t$-stat $p$-value
Clean Acc 30 7.823294 $1.255470\times 10^{-8}$ 12.387491 $4.170254\times 10^{-13}$
Adv. 30 4.348780 $1.540867\times 10^{-4}$ 5.447550 $7.317049\times 10^{-6}$
E. Adv. 30 6.595568 $3.153155\times 10^{-7}$ 9.270810 $3.572919\times 10^{-10}$
Table 9: $t$-test statistics between RFT-fix and RFT-scheduler for two perturbation magnitudes.
Model Setting $\epsilon=\nicefrac{{4}}{{255}}$ $\epsilon=\nicefrac{{8}}{{255}}$
Clean Adv. E. Adv. Clean Adv. E. Adv.
ViT fix 37.29 16.89 26.49 15.59 5.84 10.12
scheduler 70.96 22.66 46.42 63.91 11.63 34.38
SWIN fix 56.40 32.08 44.35 35.64 14.21 24.16
scheduler 79.78 39.50 60.90 72.55 22.87 47.49
CNX fix 61.14 37.59 49.93 19.82 9.61 14.33
scheduler 84.05 45.69 66.59 79.04 28.12 54.09
R50 fix 37.26 16.99 26.62 22.88 8.82 15.09
scheduler 67.76 20.13 42.65 55.57 10.48 29.84
ClipViT fix 12.73 5.74 8.85 8.59 3.13 5.56
scheduler 73.77 39.14 57.08 68.62 24.94 46.11
ClipCNX fix 24.09 14.38 18.93 13.95 7.41 10.47
scheduler 80.74 49.09 65.91 76.40 32.20 54.93
Table 10: Average per model of the clean accuracy, adversarial accuracy, and expected adversarial accuracy for $\epsilon=\nicefrac{{4}}{{255}}$ and $\epsilon=\nicefrac{{8}}{{255}}$.
Dataset Setting $\epsilon=4$ $\epsilon=8$
Clean Adv. E. Adv. Clean Adv. E. Adv.
Aircraft fix 6.37 3.47 4.66 2.58 1.77 2.14
scheduler 69.23 29.82 49.69 64.83 20.52 41.34
Caltech fix 65.42 43.00 54.27 50.37 25.67 37.51
scheduler 81.02 48.93 65.73 75.48 33.55 54.50
Cars fix 25.73 14.22 19.99 4.65 2.50 3.45
scheduler 82.43 45.33 65.29 77.25 28.45 54.16
Cub fix 47.24 23.56 35.07 19.45 5.26 11.38
scheduler 77.39 34.60 56.64 70.16 17.44 42.31
Dogs fix 46.00 18.82 31.99 19.95 5.66 11.95
scheduler 70.82 21.50 45.61 59.02 8.59 30.06
Table 11: Average per dataset of the clean accuracy, adversarial accuracy, and expected adversarial accuracy for $\epsilon=\nicefrac{{4}}{{255}}$ and $\epsilon=\nicefrac{{8}}{{255}}$.
Appendix E Additional Results on ImageNette
We provide additional results for ImageNette in Table 12
Model Setting $\epsilon=\nicefrac{{4}}{{255}}$ $\epsilon=\nicefrac{{8}}{{255}}$
Clean Adv. E. Adv. Clean Adv. E. Adv.
SWIN fix 97.15 85.07 92.22 94.19 66.17 82.43
scheduler 98.62 85.48 93.67 97.50 69.08 86.93
CNX fix 97.81 88.44 94.09 94.80 68.87 84.08
scheduler 99.29 88.23 95.36 98.27 71.32 89.03
Table 12: Imagenette results for $\epsilon=\nicefrac{{4}}{{255}}$ and $\epsilon=\nicefrac{{8}}{{255}}$.